Grade 12Physical Chemistry

The vapour pressure of pure A is 10 and at same temperature when 1 g of B is dissolved in 20 g A, its vapour pressure is reduced to 9. The molecular weight of A is 200 amu then molecular weight of B is?

Mit patel
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
Us ethe Solutions and Colliative Properties chapter formulae that is:                       
        Let PO = 10 ,Ps = 9
(P^o-P_s)/P_s =  n_B/N_A 
                                                     (10-9)/9=  □(□(1/M_B )) /□(20/200)
                                                          Therefore, MB = 90amu.